Herbert is 5 times as old as mike. If the sum of their ages is 18 years, find their ages.
Answer by checkley79(3341)

x+5x=18
6x=18
x=18/6
x=3 the age of Mike.
5*3=15 age of Herbert.
Proof:
3+15=18
18=18